### v1

# Bi-AQUA: Bilateral Control-Based Imitation Learning for Underwater Robot Arms via Lighting-Aware Action Chunking with Transformers

Abstract
Underwater robotic manipulation is fundamentally challenged by extreme lighting variations, color distortion, and reduced visibility. We introduce Bi-AQUA, the first underwater bilateral control-based imitation learning framework that integrates lighting-aware visual processing for underwater robot arms. Bi-AQUA employs a hierarchical three-level lighting adaptation mechanism: a Lighting Encoder that extracts lighting representations from RGB images without manual annotation and is implicitly supervised by the imitation objective, FiLM modulation of visual backbone features for adaptive, lighting-aware feature extraction, and an explicit lighting token added to the transformer encoder input for task-aware conditioning. Experiments on a real-world underwater pick-and-place task under diverse static and dynamic lighting conditions show that Bi-AQUA achieves robust performance and substantially outperforms a bilateral baseline without lighting modeling. Ablation studies further confirm that all three lighting-aware components are critical. This work bridges terrestrial bilateral control-based imitation learning and underwater manipulation, enabling force-sensitive autonomous operation in challenging marine environments. For additional material, please check: https://mertcookimg.github.io/bi-aqua
I Introduction
Underwater robotic manipulation is uniquely challenging because the visual appearance of the same scene can change drastically within seconds under shifts in the spectrum, intensity, and direction of underwater lighting [1]. Such variations—caused by wavelength-dependent attenuation, scattering, turbidity, and bubbles—break the visual consistency required by standard visuomotor policies, often causing catastrophic action drift even on simple pick-and-place tasks. Existing underwater image enhancement methods improve image quality [2, 3], but they do not address the fundamental control problem: closed-loop manipulation policies must adapt both perception and action generation to rapidly varying lighting.
Traditional underwater manipulation is dominated by teleoperation [4, 5], motivating recent work on AUVs, underwater robot arms, and visuomotor learning [6, 7, 8]. However, existing approaches either rely solely on vision or use unilateral control without force feedback, and none provide mechanisms to explicitly model lighting as a latent factor influencing visual observations.
Figure 1: Concept of Bi-AQUA.
In terrestrial environments, imitation learning (IL) through leader–follower teleoperation has proven highly effective for acquiring complex visuomotor skills [9]. Recent systems based on ALOHA and Action Chunking with Transformers (ACT) [10, 11], Mobile ALOHA with ACT [12], and diffusion-based visuomotor policies [13] achieve impressive real-world performance. These method yet all rely on unilateral control, as shown in Fig. 2. Without force feedback, these methods struggle in contact-rich or visually ambiguous interactions.
Bilateral control addresses this limitation by exchanging both position and force, enabling demonstrations rich in haptic and visual cues and yielding more robust generalization [14], as shown in Fig. 3. Among bilateral IL frameworks, Bi-ACT [14] plays a central role by directly extending ACT to bilateral settings, offering a strong backbone for force-sensitive visuomotor learning. Recent extensions—including low-cost bilateral data collection [15], data augmentation [16, 17], the fast bilateral teleoperation and accurate dynamics model for Bi-ACT [18], and Mamba-based motion encoders [19]—demonstrate the flexibility and effectiveness of the bilateral IL paradigm. However, all such work assumes visually stable land environments and does not address the severe lighting variability characteristic and underwater settings.
Figure 2: Unilateral Control-based Imitation Learning
A key gap therefore remains. To the best of our knowledge, there exists no imitation learning framework that (i) brings bilateral control-based imitation learning (Bi-IL) to underwater robot manipulation, nor (ii) explicitly models underwater lighting as a latent factor in force-sensitive visuomotor control. Bi-AQUA addresses both limitations simultaneously.
To close this gap, we introduce Bi-AQUA, the first underwater Bi-IL framework that explicitly models lighting at multiple levels of the visuomotor pipeline, as shown in Fig. 1. Bi-AQUA builds upon Bi-ACT–style action chunking [14, 20] and incorporates a hierarchical lighting adaptation mechanism: (1) a label-free Lighting Encoder that learns lighting representations implicitly through the imitation objective, (2) FiLM-based visual feature modulation for lighting-aware perception, and (3) a lighting token added to the transformer encoder input for lighting-adaptive action generation. Together, these components enable both perception and control to remain stable under static and dynamically changing lighting while leveraging rich force cues from bilateral demonstrations.
Our contributions are as follows:
  • Bi-AQUA is a first bilateral control–based imitation learning framework for underwater robot arms.
  • We propose a lighting-aware visuomotor policy that integrates a label-free, implicitly supervised Lighting Encoder, FiLM modulation, and a lighting token for adaptive control.
  • We demonstrated real-world gains over baseline and generalization to unseen lighting conditions, novel objects, and visual disturbances in underwater environments.
II Related Work
II-A Underwater Vision and Robot
Underwater perception is degraded by wavelength-dependent attenuation, scattering, turbidity, and non-uniform lighting [1], motivating extensive work in image enhancement: physics-based [21], learning-based [22], hybrid approaches [23, 24, 3], large datasets [25], and low-light or non-uniform illumination correction [2, 26]. While some methods target robotics [3, 2], they optimize image quality rather than manipulation performance.
Underwater manipulation research has traditionally relied on teleoperation [4, 27, 5], prompting efforts toward autonomous control of underwater vehicles and manipulators [28, 29]. Recent works explore RL and IL for navigation and manipulation [30, 6, 31] and propose systems such as AquaBot [7] or simulated benchmarks [8]. However, as summarized in the Table I, existing systems either depend solely on vision or lack integration of force feedback and lighting modeling, leaving a gap in real underwater manipulation under dynamic lighting.
Our Bi-AQUA treats lighting as a latent variable inside the visuomotor policy, learning a label-free lighting representation that conditions perception (FiLM) and sequence-level action generation (lighting token) under dynamic underwater illumination.
Figure 3: Bilateral Control-based Imitation Learning
TABLE I: Comparison of representative approaches.
Method Domain RGB Vision Underwater Bilateral Control Robot Policy Lighting Modeling
SelfLUID-Net [2] Image Enhancement
DeepSeeColor [3] Image Enhancement
RoLight [32] Manipulation
Underwater Bilateral [27, 33] Teleoperation
Ocean One [34, 35] Teleoperation
U0 [8] Manipulation/Navigation (IL:sim)
UIVNAV [6] Navigation(IL)
AquaBot [7] Manipulation (IL/RL)
Bi-IL(w/o vision) [19, 36, 37] Manipulation (IL)
ACT [10] Manipulation (IL)
Bi-ACT [14, 15, 18] Manipulation (IL)
Bi-AQUA (Ours) Manipulation (IL)
II-B Bilateral Control-based Imitation Learning
Unilateral teleoperation such as ALOHA and Mobile ALOHA [10, 11, 12] relies solely on visual feedback, and the lack of force feedback limits robustness in contact-rich manipulation.
Bilateral control-based imitation learning (Bi-IL) alleviates this limitation by exchanging both position and force between leader and follower robots, enabling demonstrations rich in interaction dynamics [36, 19, 9]. However, most Bi-IL systems use little or no visual input and instead depend almost entirely on joint and force signals. Recent work has begun integrating images into Bi-IL, with Bi-ACT extending ACT—originally developed for unilateral teleoperation [10, 11, 12]—into a Bi-IL framework [14, 15]. These methods combine force-rich demonstrations with high-dimensional perception but have only been evaluated in air with stable lighting.
Underwater environments introduce scattering, color attenuation, turbidity, and dynamic lighting, severely degrading visual observations. Consequently, proprioception-only Bi-IL becomes perceptually blind, and existing vision-based Bi-IL lacks mechanisms to explicitly model lighting, despite its major role in underwater image formation. To our knowledge, no prior work unifies Bi-IL, underwater deployment, and explicit lighting modeling within a single visuomotor policy. Bi-AQUA is the first Bi-IL framework explicitly designed for real underwater manipulation under dynamic lighting.
II-C Robust Visual Policies and Lighting Modeling
Lighting variability is one of the dominant factors degrading visuomotor policy performance, particularly in underwater environments where color attenuation, scattering, and non-uniform illumination cause rapid, appearance-altering shifts in RGB observations. Prior work has attempted to improve robustness through domain randomization [38, 39, 40, 41] and physically based relighting [32]. While effective for data augmentation, these strategies do not provide the policy with an explicit, internal representation of lighting—leaving the visuomotor pipeline to implicitly cope with color and illumination changes.
FiLM [42] offers a flexible mechanism for conditional feature modulation and has been widely used in preference-aware and multi-task manipulation [43, 44, 45]. However, existing applications typically treat lighting only implicitly—through global conditioning or augmentation—without identifying lighting as a distinct latent factor that affects both perception and downstream control.
In contrast, we explicitly model lighting as a structured internal variable. A label-free Lighting Encoder extracts lighting embeddings directly from raw RGB and is trained solely through the imitation objective, requiring no lighting labels. These embeddings modulate the visual backbone via FiLM and are injected as a dedicated lighting token into the transformer encoder for sequence-level conditioning.
Together, these components allow the visuomotor policy to adapt its perception and control to rapidly changing underwater illumination, in ways that existing domain-randomized or augmentation-only systems cannot. Bi-AQUA is therefore the first bilateral control-based imitation learning framework to incorporate explicit, hierarchical lighting modeling for real underwater manipulation.
III Bi-AQUA Method
III-A Overview
Bi-AQUA inherits Bi-ACT’s leader–follower bilateral control, leveraging both position and force information, and introduces lighting-aware visual processing at three levels: (i) a Lighting Encoder that extracts compact lighting embeddings from RGB images, (ii) FiLM-based feature-wise modulation of backbone features conditioned on the lighting embedding, and (iii) an explicit lighting token added to the transformer encoder input, which flows through the encoder to the decoder for sequence-level conditioning. This hierarchical design enables the policy to adapt to diverse and time-varying underwater lighting while preserving force-sensitive manipulation.
III-B Data Collection
Figure 4: Data collection of Bi-AQUA.
Figure 5: Overview of Bi-AQUA. Given multi-view underwater observations and follower joint states, Bi-AQUA extracts lighting-aware visual features, fuses them with proprioception, and predicts leader-side action chunks within a bilateral control loop.
Data collection follows the bilateral control paradigm of Bi-ACT, adapted to an underwater follower as shown in Fig. 4. A human operator manipulates a leader robot in air, and a follower robot executes mirrored motions underwater.
Bilateral control enforces simultaneous position tracking and force feedback between leader and follower according to
$\theta_{l}-\theta_{f}=0,\quad\tau_{l}+\tau_{f}=0,$ (1)
where $\theta$ and $\tau$ denote joint position and torques, respectively. Position is obtained from encoders, velocities by differentiation, and torques by a disturbance observer (DOB) [46] and a reaction torque observer (RTOB) [47], without dedicated force sensors.
We record multi-modal observations from the underwater follower: RGB images from $N_{c}$ cameras at frequency $f_{\text{img}}$, follower joint data (position, velocity, torque) at $f_{\text{joint}}$, and synchronized leader–follower joint states $\mathbf{l}_{t},\mathbf{f}_{t}\in\mathbb{R}^{3N}$. All streams are temporally standardized to the training frequency $f_{\text{train}}$: images are aligned by selecting the nearest frame to each training timestamp, and joint data are downsampled by taking every $\lfloor f_{\text{joint}}/f_{\text{train}}\rfloor$ sample. Images are resized to $H_{\text{img}}\times W_{\text{img}}$ via letterboxing. After preprocessing, all modalities are aligned at $f_{\text{train}}$ and stored in a unified dataset.
To ensure robustness, demonstrations are collected under diverse lighting spectra and intensities. Bilateral control allows the operator to sense contact forces, enabling safe and precise underwater manipulation during teleoperation.
III-C Learning Model
Bi-AQUA extends the Bi-ACT architecture by incorporating lighting-aware visual processing into a transformer-based action prediction model, as shown in Fig. 5. The system consists of:
  1. 1.
    a label-free lighting-aware visual encoder,
  2. 2.
    a joint state encoder,
  3. 3.
    a transformer-based action prediction module,
  4. 4.
    bilateral control integration for closed-loop execution.
III-C1 Lighting-Aware Visual Processing
Figure 6: Lighting Encoder. A convolutional path captures spatial lighting cues, while a histogram path models saturation–value statistics. The two paths are fused into a compact lighting representation without any lighting annotations.
Given multi-view observations $\mathbf{I}=\{\mathbf{I}_{i}\}_{i=1}^{N_{c}}$, our goal is to extract manipulation-relevant features while explicitly modeling lighting. We introduce three components: a Lighting Encoder, a FiLM-modulated backbone, and a lighting token.
Lighting Encoder.
As shown in Fig. 6, we design a Lighting Encoder $\mathcal{E}_{L}:\mathbb{R}^{3\times H\times W}\rightarrow\mathbb{R}^{d_{L}}$ that produces a compact lighting representation $\mathbf{v}_{L}\in\mathbb{R}^{d_{L}}$ from RGB images without lighting or color annotations, which are difficult to define in underwater settings.
The encoder adopts a dual-path architecture. A convolutional path processes the input through $L_{\text{conv}}$ convolutional layers with channels $3\to C_{1}\to C_{2}\to C_{3}$, followed by ReLU and global average pooling to yield spatial features $\mathbf{f}_{\text{conv}}\in\mathbb{R}^{C_{3}}$. A histogram path computes a 2D histogram $\mathbf{H}\in\mathbb{R}^{B_{\text{hist}}^{2}}$ over saturation and value (SV) channels and passes it through a two-layer MLP to obtain $\mathbf{f}_{\text{hist}}\in\mathbb{R}^{d_{\text{hist}}}$. The final lighting embedding is
$\mathbf{v}_{L}=\mathbf{W}_{L}[\mathbf{f}_{\text{conv}};\mathbf{f}_{\text{hist}}]\in\mathbb{R}^{d_{L}},$ (2)
with $\mathbf{W}_{L}\in\mathbb{R}^{d_{L}\times(C_{3}+d_{\text{hist}})}$ and concatenation $\cdot$$\cdot$]. For multiple cameras, we use a shared encoder and average across views:
$\mathbf{v}_{L}=\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\mathcal{E}_{L}(\mathbf{I}_{i}).$ (3)
FiLM-modulated visual backbone.
Following the formulation of FiLM [42], we modulate each convolutional feature map via a feature-wise affine transformation whose parameters are generated from the lighting embedding. For a feature map $\mathbf{F}_{\ell}\in\mathbb{R}^{B\times C_{\ell}\times H_{\ell}\times W_{\ell}}$ at layer $\ell$, the FiLM operation is
$\operatorname{FiLM}(\mathbf{F}_{\ell}\mid\boldsymbol{\gamma}_{\ell},\boldsymbol{\beta}_{\ell})=(1+\boldsymbol{\gamma}_{\ell})\odot\mathbf{F}_{\ell}+\boldsymbol{\beta}_{\ell},$ (4)
where $\boldsymbol{\gamma}_{\ell},\boldsymbol{\beta}_{\ell}\in\mathbb{R}^{C_{\ell}}$ are per-channel scaling and shifting coefficients, broadcast spatially over $(H_{\ell},W_{\ell})$, and $\odot$ denotes element-wise multiplication. This formulation preserves the identity mapping when $\boldsymbol{\gamma}_{\ell}=\mathbf{0}$ and $\boldsymbol{\beta}_{\ell}=\mathbf{0}$, facilitating stable training initialization.
Both parameters are produced by learned functions of the conditioning input—here, the lighting vector $\mathbf{v}_{L}$:
$\boldsymbol{\gamma}_{\ell}=f_{\ell}(\mathbf{v}_{L}),\qquad\boldsymbol{\beta}_{\ell}=h_{\ell}(\mathbf{v}_{L}),$ (5)
where $f_{\ell}$ and $h_{\ell}$ are implemented as linear projections
$f_{\ell}(\mathbf{v}_{L})=\mathbf{W}_{\gamma}^{(\ell)}\mathbf{v}_{L},\qquad h_{\ell}(\mathbf{v}_{L})=\mathbf{W}_{\beta}^{(\ell)}\mathbf{v}_{L},$ (6)
with $\mathbf{W}_{\gamma}^{(\ell)},\mathbf{W}_{\beta}^{(\ell)}\in\mathbb{R}^{C_{\ell}\times d_{L}}$. In our implementation, FiLM is applied to the final ResNet layer. The FiLM-modulated features are then flattened into tokens, augmented with 2D sinusoidal positional embeddings, and concatenated across camera views to form the visual token sequence $\mathbf{F}_{\text{vis}}\in\mathbb{R}^{N_{\text{tokens}}\times d_{h}}$ used by the transformer decoder.
III-C2 Training Objective
Bi-AQUA employs a Conditional Variational Autoencoder (CVAE) with a transformer backbone, following Bi-ACT while augmenting it with lighting-aware conditioning. Given follower state $\mathbf{f}_{t}$ and visual observations $\mathbf{I}_{t}$, the policy predicts a chunk of leader actions $\hat{\mathbf{l}}_{t:t+k}$. The lighting token is added to the transformer encoder input alongside latent and proprioceptive tokens. The encoder processes these tokens together, and the resulting memory (which includes the lighting information) is then accessed by the transformer decoder via cross-attention, enabling the decoder to adapt action generation to the current underwater lighting.
Figure 7: Inference pipeline of Bi-AQUA.
We train Bi-AQUA end-to-end using behavior cloning with a CVAE-style latent space:
$\mathcal{L}=\mathcal{L}_{\text{action}}+\lambda_{\text{KL}}\mathcal{D}_{\text{KL}},$ (7)
where $\mathcal{L}_{\text{action}}$ is an L1 loss between predicted and ground-truth leader actions, and $\mathcal{D}_{\text{KL}}$ is the KL divergence between the learned latent distribution and a standard normal prior, weighted by $\lambda_{\text{KL}}$. The Lighting Encoder, FiLM layers, and transformer modules are optimized jointly.
III-D Inference
At test time, Bi-AQUA runs in closed loop with the bilateral controller, as shown in Fig. 7. At each inference step, the model receives current multi-view images and follower joint state, computes $\mathbf{v}_{L}$ with the Lighting Encoder, applies FiLM modulation to the backbone, samples a latent action code from the prior, and uses the transformer decoder to generate an action chunk for the leader robot.
The predicted leader actions are converted into follower commands by the bilateral controller and executed at control frequency $f_{\text{control}}$. Since the policy operates at a lower inference rate $f_{\text{inference}}$, we maintain an action chunk buffer of length $k$ and output actions at $f_{\text{train}}$ between inference calls, ensuring smooth and responsive motion.
The combination of Lighting Encoder, FiLM-based visual modulation, and a lighting token yields a lighting-aware visuomotor policy that maintains reliable underwater manipulation even under challenging and dynamically changing lighting.
IV Experiments
IV-A Experimental Setup
IV-A1 Hardware Configuration
Figure 8: Experimental Environments
We evaluate Bi-AQUA on a bilateral leader–follower system, as shown in Fig. 8. Each arm has three joints (J1–J3) and a gripper (J4), driven by Dynamixel XW540-T260 (IP68) servo motors. The follower arm is mounted on a 200 mm pedestal inside a water tank of $900\times 450\times 450$ $m$.
A multi-color RGBW LED aquarium light (FEDOUR RD10-400RGBW, 38 cm, 7 W, 21 LEDs, IP68) is installed beneath the follower pedestal and provides eight reproducible lighting conditions. All lighting modes used in our experiments (red, blue, green, white, rgbw, cyan, purple, changing) correspond to the manufacturer’s 100% brightness presets. For dynamic changing mode, we use the device’s built-in “M5” mode, which automatically cycles through all colors at 2 s intervals. This setup ensures consistent spectral output across trials while allowing us to systematically evaluate robustness under both static and rapidly changing underwater lighting.
Bi-AQUA receives RGB observations from two cameras(TIER IV C1 120): a gripper-mounted camera for close-up views and a top camera capturing the global workspace, including the basket.
IV-A2 Task and Evaluation Protocol
Figure 9: Teleoperation data collection for the pick-and-place task under five lighting modes (e.g., white).
The task is a pick-and-place designed to stress lighting robustness. The follower must grasp a block on the right side of the tank and place it into a basket on the opposite side, passing through four stages: (#1) initial pose, (#2) pick up, (#3) transport, (#4) place.
For teleoperated data collection, as shown in Fig. 9, we execute the water task twice under each of five lighting conditions (red, blue, green, white, rgbw), yielding 10 demonstrations. The primary training object is a polyurethane block, and the operator can adapt motions to visibility while bilateral control provides consistent force feedback.
For autonomous evaluation, we test all eight lighting conditions, including the dynamic changing mode and two unseen colors (cyan, purple). To assess appearance generalization, we additionally evaluate: (a) a black rubber block, (b) a blue sponge, and (c) bubbles around the polyurethane block as shown in Fig. 10. Each configuration is evaluated over 5 autonomous rollouts and considered successful if the object ends inside the basket without being dropped.
Figure 10: Example task for generalization to novel objects and disturbances
Figure 11: Overview of collected datasets
IV-B Dataset and Training Configuration
IV-B1 Demonstration Collection and Preprocessing
We collect 10 bilateral teleoperation episodes (two per training lighting condition). All demonstrations execute the same pick-and-place task; the operator follows an identical intended motion pattern across episodes so that the primary variation comes from lighting rather than from changes in the manipulation strategy. The policy is trained and evaluated solely from follower-side observations; leader-side joint states serve as action labels.
Two RGB cameras record at 30 Hz with $1920\times 1280$ resolution: (1) a top camera above the water tank, (2) a gripper camera near the follower gripper. Images are resized to $480\times 320$. Joint position, velocity, and torque for $N=4$ joints (J1–J4) are recorded at 1,000 Hz from both arms. Leader joint states are used as actions; follower joint states are part of the observation. For training, joint data are downsampled to 100 Hz, and images are temporally aligned to 100 Hz by selecting the frame closest to each 10 ms interval. Each time step in the final dataset therefore contains (a) two RGB views (top and gripper), (b) follower joint position, velocity, and torque, and the corresponding leader joint states. Across the 10 demonstrations, this results in 15,402 time steps, corresponding to 30,804 RGB frames in total (two views per time step) and 15,402 synchronized joint‑state vectors; each contains 24 values (leader+follower, 4 joints × position/velocity/torque), totaling 369,648 values. The overview of collected datasets is shown in Fig. 11
Although each episode is collected under a known lighting mode (e.g., red, blue, green), we do not provide any lighting IDs or manual labels to the policy. The Lighting Encoder is trained end-to-end from the control objective and learns a latent representation of lighting purely from visual input.
IV-B2 Network Architecture
TABLE II: Ablation of lighting-aware components in Bi-AQUA.
Method Lighting Encoder Lighting Token FiLM
Bi-ACT (Baseline)
Bi-ACT+LE-Token
Bi-ACT+LE-FiLM
Bi-AQUA (Ours)
All model variants share the same Bi-ACT-style visuomotor backbone; the ablations in Table II differ only in the presence or absence of three lighting-aware components. Concretely, the architecture is controlled by three binary switches: (a) a Lighting Encoder that produces a 64-D lighting embedding, (b) FiLM modulation applied to the visual backbone, and (c) a lighting token added to the transformer encoder input. Table II summarizes which of these components is enabled for each variant.
IV-B3 Training Configuration
All models are trained offline using the same 10 demonstrations, preprocessing pipeline, and optimization hyperparameters to ensure a fair comparison across the rows of Table II. We optimize a Bi-ACT-style action chunks with an additional KL regularization term for the CVAE latent space, using AdamW for both the ResNet-18 backbone and the remaining modules and KL weight $\lambda_{\text{KL}}=1.0$.
At deployment, the trained policies run in real time on the physical robot using an NVIDIA RTX 3060 Ti GPU (8 GB VRAM).
Figure 12: Autonomous execution of Bi-AQUA under lighting conditions.
TABLE III: Task success rate (%) under different lighting conditions.
Method Trained Lighting Untrained Lighting
red blue green white rgbw cyan purple changing
Bi-ACT (Baseline) 20 0 0 100 0 0 0 0
Bi-ACT+LE-Token 0 0 0 0 0 0 0 0
Bi-ACT+LE-FiLM 80 100 100 100 100 60 100 20
Bi-AQUA (Ours) 100 80 100 100 100 100 100 100
Bi-AQUA generalization to novel objects and disturbances
Black Block 60 0 60 100 100 100 100 60
Sponge 100 20 100 20 0 80 20 40
Bubble 20 80 80 100 80 0 100 40
IV-C Experimental Results
IV-C1 Lighting Robustness and Ablation
Table III summarizes task success rates across all eight lighting conditions, including three unseen at training time (cyan, purple, changing). Fig. 12 shows the autonomous execution of Bi-AQUA. Bi-AQUA achieves consistently high robustness, attaining 100% success in seven of the eight lighting conditions and maintaining 80% success even under blue light, which is the most challenging condition due to severe wavelength-dependent attenuation. Crucially, Bi‑AQUA achieves 100% and clearly outperforms all baselines in the dynamic changing mode, where lighting cycles through all spectra every 2 sec. This setting induces rapid, intra-horizon appearance shifts that invalidate the stationarity assumptions of standard visuomotor policies.
In contrast, the lighting-agnostic Bi-ACT baseline succeeds only under white (100%) and partially under red (20%), and fails completely in the remaining six conditions despite using the same demonstrations. This sharp performance gap demonstrates that underwater visuomotor control fundamentally requires explicit lighting modeling rather than relying on passive robustness of the visual backbone.
The ablations in Table II isolate the role of each lighting-aware component. Bi-ACT+LE-Token, which uses a Lighting Encoder to extract lighting embeddings and adds a lighting token to the transformer encoder input (conditioning the decoder via memory) without modifying the visual features via FiLM, fails in all lighting conditions. This indicates that high-level conditioning alone is insufficient for correcting color shifts and contrast changes introduced by underwater propagation.
Bi-ACT+LE-FiLM, which uses a Lighting Encoder to extract lighting embeddings and applies FiLM modulation to the final ResNet layer (layer4) but omits the lighting token, performs substantially better, achieving up to 100% success in five static or unseen lighting conditions. However, its success drops sharply to 20% in the dynamic changing mode, suggesting that FiLM alone cannot track rapid temporal lighting transitions.
Only the full Bi-AQUA model—combining Lighting Encoder, FiLM modulation, and a sequence-level lighting token—maintains uniformly high performance across all lighting conditions. This confirms that hierarchical lighting reasoning, spanning both feature modulation and sequence-level conditioning, is essential for underwater manipulation under severe lighting variability.
Fig. 13 further analyzes task execution time. Averaged over all lighting conditions, Bi-ACT takes $20.17$ sec per successful rollout, Bi-ACT+LE-FiLM $19.61$ sec, Bi-AQUA $15.73$ sec, and human teleoperation (Train) $15.39$ sec. Although Bi-ACT+LE-FiLM achieves high success in many static lighting settings, it consistently completes the task several seconds slower than Bi-AQUA and the teleoperator, reflecting cautious, stop-and-go behavior under difficult colors. Bi-AQUA, by contrast, closely matches the teleoperation reference in both mean time and variance, suggesting that explicit lighting modeling not only increases success rates but also preserves efficient, human-like trajectories instead of sacrificing speed for robustness. The per-lighting breakdown in Fig. 13 shows that this advantage holds across all lighting modes, including the dynamic changing condition, with Bi-AQUA maintaining consistently short and stable execution times.
Figure 13: Execution time profiles across the task stages under different lighting conditions.
IV-C2 Generalization to Novel Objects and Disturbances
To evaluate whether Bi-AQUA overfits to the polyurethane block used during training, we assess three additional evaluation settings: a black rubber block, a blue sponge, and bubbles surrounding the manipulated object. The results in Table III demonstrate that Bi-AQUA generalizes well beyond the training distribution despite having only ten demonstrations.
For the black block, Bi-AQUA achieves success rates between 60% and 100%, with perfect performance under white, rgbw, cyan, and purple lighting. This indicates robustness to drastic changes in albedo and object contrast.
For the blue sponge, which introduces significant geometric, textural, and chromatic variation, performance varies more widely (0–100%). Even under these severe appearance shifts, Bi-AQUA maintains non-trivial success rates across multiple lighting modes, achieving 80% under cyan and 40% in the dynamic changing mode.
The bubble condition is particularly challenging from a perception standpoint. Because the top camera observes the scene through the water surface from above, air injection generates surface waves that induce strong refraction and temporal warping, severely degrading the global view of the tank. At the same time, bubbles rise around the gripper and the object, intermittently occluding the local workspace and creating high-frequency specular highlights and false edges in the gripper view. Despite these compounded distortions, Bi-AQUA still achieves up to 100% success in several lighting conditions, as shown in Table III. This indicates substantial resilience to simultaneous global (water-surface) and local (near-gripper) disturbances.
Overall, these results show that Bi-AQUA not only provides strong lighting robustness but also extends to unseen objects and disturbance conditions without any additional data or fine-tuning. The remaining failure cases point to the inherent difficulty of underwater visual degradation and motivate future work on improved relighting models, physics-based augmentation, and multimodal sensing.
V Conclusion
We introduced Bi-AQUA, a bilateral control–based imitation learning framework for underwater manipulation that explicitly models lighting at multiple levels. Through a label-free Lighting Encoder implicitly supervised by the imitation objective, FiLM-based conditioning of visual features, and a sequence-level lighting token added to the transformer encoder input, Bi-AQUA adapts its visuomotor behavior to changing lighting while preserving the force-sensitive advantages of bilateral teleoperation. Ablation and generalization experiments show that the lighting-aware components provide complementary gains and that the learned lighting representation transfers to novel objects and visual disturbances.
Limitations and future work. Our evaluation is restricted to a single task, robot platform, and tank environment with a finite set of lighting modes, and performance still degrades under extreme combinations of lighting and object appearance. We also do not yet consider variation in water quality, background clutter, or large-scale field deployments. Extending Bi-AQUA to richer skill repertoires, more realistic subsea settings, and integration with other adaptation mechanisms (e.g., domain randomization or online adaptation) are important directions for future work.